# Characteristic Guidance: Non-linear Correction for DDPM at Large Guidance Scale

### v1

@@ 4.2 Theoretical Foundation
The characteristic guidance (5) is derived from the Fokker Planck equation (Appendix C) governing the DDPM under the harmonic ansatz 4.1.
Before we proceed, let’s define
$\boldsymbol{\epsilon}(\mathbf{x},t)=\operatorname*{argmin}_{\boldsymbol{\epsilon}_{\theta}}L(\boldsymbol{\epsilon}_{\theta})$
as the optimal solution of the DDPM’s de-noising objective (2) with infinitesimal diffusion time steps. We propose the Harmonic ansatz stating that such optimal solution is harmonic:
Ansatz 4.1.
@@
$\nabla_{\mathbf{x}}^{2}\boldsymbol{\epsilon}(\mathbf{x},t)=0.$
This ansatz is inspired by and holds exactly when the distribution of original images $p(\mathbf{x})$ is a Gaussian distribution, and we will show that it works for other cases as well. Although not universally applicable, the harmonic ansatz is a less restrictive assumption compared to the linear additivity in classifier-free guidance, which ignores all non-linear effects.
The characteristic guidance is a direct consequence of the following lemma
@@
Let $\boldsymbol{\epsilon}(\mathbf{x},t)$, $\boldsymbol{\epsilon}_{1}(\mathbf{x},t)$, and $\boldsymbol{\epsilon}_{2}(\mathbf{x},t)$ be optimal solutions of three distinct DDPMs satisfying the Harmonic Ansatz 4.1. Moreover, their initial condition at $t=0$ satisfies
$\boldsymbol{\epsilon}(\mathbf{x},0)=(1+\omega)\boldsymbol{\epsilon}_{1}(\mathbf{x},0)-\omega\boldsymbol{\epsilon}_{2}(\mathbf{x},0),$
where $\omega$ is a constant. Then, we have the relation
$\boldsymbol{\epsilon}(\mathbf{x},t)=(1+\omega)\boldsymbol{\epsilon}_{1}(\mathbf{x}+\omega\Delta\mathbf{x},t)-\omega\boldsymbol{\epsilon}_{2}(\mathbf{x}+(1+\omega)\Delta\mathbf{x},t),$
@@
$\Delta\mathbf{x}=\left(\boldsymbol{\epsilon}_{2}(\mathbf{x}+(1+\omega)\Delta\mathbf{x},t)-\boldsymbol{\epsilon}_{1}(\mathbf{x}+\omega\Delta\mathbf{x},t)\right)\sigma(t),$
in which $\sigma(t)=\sqrt{1-e^{-t}}$ for $t\in[0,T]$ be the time of forward diffusion process.
Replacing $\boldsymbol{\epsilon}$, $\boldsymbol{\epsilon}_{1}$, and $\boldsymbol{\epsilon}_{2}$ with $\boldsymbol{\epsilon}_{CH}(\mathbf{x}|\mathbf{c},t_{i},\omega)$, $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}|\mathbf{c},t_{i})$, and $\boldsymbol{\epsilon}_{\theta}(\mathbf{x},t_{i})$ gives us the characteristic guidance. The detailed argument for lemma 4.2 and the connection between $\sigma(t)$ and $\sigma_{i}$ are established in Appendix E. The characteristic guidance applies the classifier-free guidance only at the initial condition $t_{0}=0$, and then solves the Fokker-Planck equation under the harmonic ansatz for the rest of the diffusion process. This allows us to use the classifier-free guidance when it is most accurate and avoid its drawbacks later on.
@@ 5 Experiments
5.1 Conditional Gaussian: Where the Harmonic Ansatz Holds Exactly
We compare the performance of classifier-free guidance and characteristic guidance on sampling from conditional 2D Gaussian distributions, a scenario where the harmonic ansatz is exact.
@@ 5.1 Conditional Gaussian: Where the Harmonic Ansatz Holds Exactly
Figure 2: Samples and the KL divergence from characteristic guidance (CH) and classifier free guidance (CF) guided DDPM modeling conditional Gaussian distribution. Samples are drawn with various sample methods: SDE, probabilistic ODE, DDIM and DPM++2M. The contours corresponds to the theoretical reference distribution of the guided DDPM (3).
Two DDPMs are analytically defined (Appendix F) for two Gaussian distribution: the conditional distribution $p(x_{1},x_{2}|c_{1},c_{2})$ and the unconditional distribution $p(x_{1},x_{2})$. Both of them are 2D Gaussian distributions
$\begin{split}p(\mathbf{x}|\mathbf{c})&=\mathcal{N}\left(x_{1},x_{2}|(c_{1},c_{2})^{T},I\right)\\ p(\mathbf{x})&=\mathcal{N}\left(x_{1},x_{2}|(0,0)^{T},5I\right)\end{split}$ (7)
The guided diffusion model (27) aims to sample from the distribution (3), which is another 2D Gaussian distribution for $x_{1},x_{2}$ whose contour is plotted in Fig.2. The harmonic ansatz holds exactly for Gaussian distributions because their score functions are harmonic, so as the corresponding DDPMs modeling them.
Samples from the classifier-free guidance and characteristic guidance at $c_{1}=-5,c_{2}=5,\omega=4$ are shown in Fig.2. More results could be found in Fig.8. The projection operator $\mathbf{P}$ for characteristic guidance is set to be identity. Four different sampling methods are tested: the SDE (23) (1000 steps), probabilistic ODE (Song et al., 2020b) (1000 steps), DDIM (Song et al., 2020a) (20 steps), and DPM++2M (Lu et al., 2022) (20 steps). These samples are compared with the theoretical reference distribution (LABEL:exp_Gaussian_guided_diffusion_analytical) using Kullback–Leibler(KL) divergence, which is calculated by assuming the samples are also Gaussian distributed.
Fig.2 shows that characteristic guidance achieves better KL divergence than classifier-free guidance for every sampling methods. It worth noting that classifier-free guidance suffers from severe bias and catastrophic loss of diversity when ODE based sampling methods ( ODE, DDIM, and DPM++2M ) are used. Contrarily, characteristic guidance always yields correct sampling with KL divergence better than classifier-free guidance’s best SDE samples. In conclusion, characteristic guidance is capable of reducing bias, increase diversity, and eliminating the performance gap between ODE and SDE sampling methods for analytically defined DDPMs when the harmonic ansatz holds exactly.
@@
Figure 3: Comparison between characteristic guidance (CH) and classifier free guidance (CF) guided DDPM on modeling mixture of Gaussian distribution. Samples are drawn with various sample methods: SDE, probabilistic ODE, DDIM and DPM++2M. The contours corresponds to the theoretical reference distribution of the guided DDPM (3). Samples from characteristic guidance shows better KL divergence than those from classifier-free guidance.
5.2 Mixture of Gaussian: Where the Harmonic Ansatz Holds Roughly
@@
We compare the performance of classifier-free guidance and characteristic guidance on sampling from Mixture of Gaussian distributions. In this case, the harmonic ansatz is a good approximation for most of the region except where the components overlap significantly.
We train conditional and unconditional DDPM to learn two distributions: The conditional distribution $p(\mathbf{x}|\mathbf{c})$
$p(\mathbf{x}|\mathbf{c})=\prod_{i=0}^{2}\mathcal{N}\left(\mathbf{x}|\boldsymbol{\mu}_{i},I\right)^{c_{i}},$ (8)
where $\mathbf{c}=(c_{0},c_{1},c_{2})^{T}$ is a three dimensional one-hot vector, $\boldsymbol{\mu}_{0}=(-1,-1/\sqrt{3})^{T}$, $\boldsymbol{\mu}_{1}=(1,-1/\sqrt{3})^{T}$, $\boldsymbol{\mu}_{2}=(0,\sqrt{3}-1/\sqrt{3})^{T}$, and the unconditional distribution $p(\mathbf{x})$
$p(\mathbf{x})=\frac{1}{3}\sum_{\{\mathbf{e}_{i},i=0,1,2\}}p(\mathbf{x}|\mathbf{e}_{i})$ (9)
where $\{\mathbf{e}_{i},i=0,1,2\}$ represents all one-hot vectors in three dimension space. The guided diffusion model is again aims to sample from the distribution (3), whose contour is plotted in Fig.3.
@@
Samples from the classifier-free guidance and characteristic guidance at $\omega=6$ are shown in Fig.3. For characteristic guidance, the projection operator $\mathbf{P}$ is set to be identity. More results could be found in Fig.52. Four different sampling methods are tested: the SDE (1000 steps), probabilistic ODE (1000 steps), DDIM (20 steps), and DPM++2M (20 steps). Samples are compared with the theoretical reference distribution (3) using Kullback–Leibler(KL) divergence, which is calculated by assuming the sample distribution is another mixture of Gaussian with each components are Gaussian distributed.
@@
Fig.3 shows that the classifier-free guidance again suffers from bias and significantly loss of diversity when ODE based sampling methods ( ODE, DDIM, and DPM++2M ) are used. Contrarily, samples from characteristic guidance has less bias, better diversity, and smaller performance gap between ODE and SDE. However, these samples has less clear boundaries between components than the theoretical reference contour. This is because the harmonic ansatz is not valid in the overlap regions. Besides, minor artifacts appears because iterating on trained neural networks $\boldsymbol{\epsilon}$ brings in approximation errors. Nonetheless, characteristic guidance is still capable of reducing bias, improve diversity, and reduce the performance gap between ODE and SDE sampling methods when the harmonic ansatz holds roughly.
@@ 5.3 Cooling of the Magnet: Where the Harmonic Ansatz Not Supposed to Hold
Guidance can be interpreted as a cooling process in physics with the guidance scale $\omega$ representing the temperature drop. We demonstrate that guidance can simulate the cooling of a magnet around the Curie temperature, at which a paramagnetic material gains permanent magnetism through phase transition. Such transition is a consequence of 3rd order terms in score function, where the harmonic ansatz does not hold at all.
Figure 4: Comparison between characteristic guidance (CH) and classifier free guidance (CF) guided DDPM on simulating magnet cooling. The images in the upper right depict samples from DDPMs, while the histograms illustrate the distribution of samples’ mean value (magnetization) across different temperatures. Samples are drawn with various sample methods: SDE (see Appendix H), probabilistic ODE, DDIM and DPM++2M. The contours corresponds to the theoretical reference distribution of magnetization. The characteristic guidance has better NLL and is more capable in capturing peak separation of sample magnetization.
Figure 5: Comparative visualization of aircraft images generated from DDPM using Classifier Free Guidance (CF) versus Characteristic Guidance (CH) without cherry-picking. With increasing guidance scale ($\omega$), the images guided by CF tend to have dull or even white backgrounds. However, the images guided by CH have more vivid and realistic backgrounds with sky and clouds.
@@
Thermal movements of molecules in our magnet gradually change our magnet’s magnetization $\phi$ from one configuration to another. If we keep track of the field $\phi$ every second for a long duration, statistical physics says that the probability of observing a particular configuration of $\phi$ is proportional to the Boltzmann distribution
@@
where $\beta H(\phi;T)$ is Hamiltonian at temperature $T$, assigning an energy to each of possible $\phi$. At temperature near the Curie temperature $T_{c}$, a Landau-Ginzburg model (Kardar, 2007) of magnet uses a Hamiltonian (54) (More details are described in Appendix H) with 4-th order terms. Such terms are the key to characterise phase transition of magnet, but its gradient (score function) is a third order term that does not respect the harmonic ansatz.
@@
where $\omega>0$ is the guidance scale, $T_{0}>T_{1}$ are two distinct temperature. This means if we know the distribution of our magnet at two distinct temperatures $T_{0}$ and $T_{1}$, we know its distribution at any temperature cooler than $T_{0}$ by adjusting the guidance scale $\omega$.
@@
To simulate the cooling of our magnet with the Curie temperature set to $T_{c}=200$, we train a conditional DDPM on $\phi$ at two distinct temperatures $p(\phi|T_{0}=201)$ and $p(\phi|T_{1}=200)$ with data generated by the Metropolis-Hastings algorithm. Then we sample from the DDPM at lower temperatures by adjusting the guidance scale. Four different sampling methods are tested: the SDE (1000 steps), probabilistic ODE (1000 steps), DDIM (20 steps), and DPM++2M (20 steps). The projection operator $\mathbf{P}$ for characteristic guidance is set to be the channal-wise mean. Samples are generated again in four different ways: the SDE, probabilistic ODE, DDIM, and DPM++2M. The histogram of the mean value of the magnetization $\phi$ field values are plotted in Fig.4. More results are plotted in Fig.10.
Fig.4 demonstrates the phase transition at the Curie temperature $T_{c}=200$. Above the Curie temperature, the histogram of the mean magnetization has one peak centered at 0. This means the average magnetization of our magnet over a long time is zero, although thermal movements may lead to a non-zero net magnetization at a certain instance. Below the Curie temperature, the histogram of the mean magnetization has two peaks with non-zero centers. In this case, jumping from one peak to another becomes difficult because thermal movements of molecules can only slightly change the mean magnetization. This leads to a non-zero average magnetization over a long time, or say, a permanent magnet.
@@
Characteristic guidance outperforms classifier-free guidance in sample quality below the Curie temperature, showcasing better negative log-likelihood (NLL) and effectively capturing the phase transition with distinct peaks in magnetization of samples. In contrast, classifier-free guidance struggles to produce distinct peaks. Despite its biased peak positions compared with theoretical expectations, characteristic guidance still demonstrates superior performance in modeling phase transitions, where the harmonic ansatz does not hold.
5.4 Cifar 10: Natural Image Generation

### v5

@@ 5.2 Theoretical Foundation
This section derives the characteristic guidance (10)-(11) by solving the FP equation (4) under the Harmonic ansatz. We propose the Harmonic ansatz assuming that the optimal DDPM solution $\epsilon(\mathbf{x},t)$ of (3) is harmonic:
Ansatz 5.1.
@@
$\nabla_{\mathbf{x}}^{2}\boldsymbol{\epsilon}(\mathbf{x},t)=0.$ (13)
This ansatz is inspired by and holds exactly when the distribution of original images $p(\mathbf{x})$ is a Gaussian distribution, and we will show that it works for other cases as well in experiments 6.1 and 6.2.
Characteristic guidance adopts the same formulation as classifier-free guidance at $t=0$
$\boldsymbol{\epsilon}_{CH}(\mathbf{x}_{0}|\mathbf{c},0,\omega)=(1+\omega)\ \boldsymbol{\epsilon}(\mathbf{x}_{0}|\mathbf{c},0)-\omega\ \boldsymbol{\epsilon}(\mathbf{x}_{0},0),$ (14)
in which parameters $\theta$ are omitted since we are considering optimal solutions with infinitesimal time steps. This guarantees that $\boldsymbol{\epsilon}_{CH}$ samples from $p(\mathbf{x}|\mathbf{c},\omega)$, which aligns with the target of classifier-free guidance.
Next, our aim is to construct the unknown $\boldsymbol{\epsilon}_{CH}$ for any time $t>0$ using the known $\boldsymbol{\epsilon}(\mathbf{x}|\mathbf{c},t)$ and $\boldsymbol{\epsilon}(\mathbf{x},t)$, by solving the Fokker-Planck equation (4) analytically.
The harmonic ansatz 5.1 simplifies the Fokker-Planck equation (4) into a quasi-linear equation, which can be solved by the method of characteristics in Sec.3.4. Assuming $\boldsymbol{\epsilon}_{CH}(\mathbf{x}|\mathbf{c},t,\omega)$, $\boldsymbol{\epsilon}(\mathbf{x}|\mathbf{c},t)$, and $\boldsymbol{\epsilon}(\mathbf{x},t)$ are three solutions of (4), there exist three distinct characteristic lines $\mathbf{x}(t)$, $\mathbf{x}^{(1)}(t)$, and $\mathbf{x}^{(2)}(t)$ passing through the point $\mathbf{x}_{0}$ at $t=0$, satisfying
$\begin{split}\boldsymbol{\epsilon}_{CH}(\mathbf{x}_{0}|\mathbf{c},0,\omega)&=\mathbf{F}_{t}^{-1}\left[\boldsymbol{\epsilon}_{CH}(\mathbf{x}_{t}|\mathbf{c},t,\omega)\right],\\ \boldsymbol{\epsilon}(\mathbf{x}_{0}|\mathbf{c},0)&=\mathbf{F}_{t}^{-1}\left[\boldsymbol{\epsilon}(\mathbf{x}_{t}^{(1)}|\mathbf{c},t)\right],\\ \boldsymbol{\epsilon}(\mathbf{x}_{0},0)&=\mathbf{F}_{t}^{-1}\left[\boldsymbol{\epsilon}(\mathbf{x}_{t}^{(2)},t)\right],\end{split}$ (15)
in which the function $\mathbf{F}_{t}$ for characteristic lines of (4) is a linear function. Consequently, these characteristic lines and the initial condition (14) yield an elegant analytical solution of $\boldsymbol{\epsilon}_{CH}$:
$\boldsymbol{\epsilon}_{CH}(\mathbf{x}_{t}|\mathbf{c},t,\omega)=(1+\omega)\ \boldsymbol{\epsilon}(\mathbf{x}_{t}^{(1)}|\mathbf{c},t)-\omega\ \boldsymbol{\epsilon}(\mathbf{x}_{t}^{(2)},t).$ (16)
This result establishes the possibility of constructing $\boldsymbol{\epsilon}_{CH}$ using $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}|\mathbf{c},t_{i})$ and $\boldsymbol{\epsilon}_{\theta}(\mathbf{x},t_{i})$. Its precise form is described in the following lemma:
@@
Let $\boldsymbol{\epsilon}(\mathbf{x},t)$, $\boldsymbol{\epsilon}_{1}(\mathbf{x},t)$, and $\boldsymbol{\epsilon}_{2}(\mathbf{x},t)$ be three distinct solutions of the FP equation (4), satisfying the Harmonic Ansatz 5.1. Moreover, their initial condition satisfies
$\boldsymbol{\epsilon}(\mathbf{x},0)=(1+\omega)\boldsymbol{\epsilon}_{1}(\mathbf{x},0)-\omega\boldsymbol{\epsilon}_{2}(\mathbf{x},0),$ (17)
Then, we have the relation
$\boldsymbol{\epsilon}(\mathbf{x},t)=(1+\omega)\boldsymbol{\epsilon}_{1}(\mathbf{x}+\omega\Delta\mathbf{x},t)-\omega\boldsymbol{\epsilon}_{2}(\mathbf{x}+(1+\omega)\Delta\mathbf{x},t),$ (18)
@@
$\Delta\mathbf{x}=\left(\boldsymbol{\epsilon}_{2}(\mathbf{x}+(1+\omega)\Delta\mathbf{x},t)-\boldsymbol{\epsilon}_{1}(\mathbf{x}+\omega\Delta\mathbf{x},t)\right)\sigma(t),$ (19)
in which $\sigma(t)=\sqrt{1-e^{-t}}$.
The proof of lemma 5.2 is established in Appendix D. Replacing the continuous functions $\sigma(t)$, $\boldsymbol{\epsilon}$, $\boldsymbol{\epsilon}_{1}$, and $\boldsymbol{\epsilon}_{2}$ with their discretized version $\sigma_{i}$, $\boldsymbol{\epsilon}_{CH}(\mathbf{x}|\mathbf{c},t_{i},\omega)$, $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}|\mathbf{c},t_{i})$, and $\boldsymbol{\epsilon}_{\theta}(\mathbf{x},t_{i})$ gives us the characteristic guidance. Besides, a direct consequence of lemma 5.2 is the following Theorem
Theorem 5.3.
The characteristic guidance has no mixing error when the Harmonic ansatz applies and diffusion time steps are infinitesimal:
$\mathbf{e}_{m}(\boldsymbol{\epsilon}_{CH},\mathbf{x},t)=0.$ (20)
In summary, $\boldsymbol{\epsilon}_{CH}$ is an exact solution to the FP equation (4) with initial condition (14) under the harmonic ansatz, eliminating the mixing error caused by non-linearity of (4).
@@ 6 Experiments
6.1 Gaussians: Where the Harmonic Ansatz Holds
@@
Figure 2: Samples and the KL divergence from characteristic guidance (CH) and classifier free guidance (CF) guided DDPM modeling conditional Gaussian distribution. The contours corresponds to the theoretical reference (ground truth) distribution of the guided DDPM (5).
@@
Figure 3: Comparison between CH (ours) and CF guided DDPM on modeling mixture of Gaussian distribution. The contours corresponds to the theoretical reference distribution of the guided DDPM (5). Samples from characteristic guidance shows better KL divergence than those from classifier-free guidance.
@@
This section demonstrates the theoretical advantage of characteristic guidance on Gaussian models, where we can compare samples against analytical solutions. We conducted experiments to compare classifier-free guidance and characteristic guidance under two scenarios: sampling from (1) conditional 2D Gaussian distributions, where the harmonic ansatz is exact; (2) mixture of Gaussian distributions, where the harmonic ansatz holds for most of the region except where the components overlap significantly. The guided diffusion model (39) aims to sample from the distribution (5) whose contour is plotted in Fig.2 and Fig.3.
In the conditional Gaussian scenario, we analytically set two DDPMs for the conditional distribution $p(\mathbf{x}|\mathbf{c})$ and the unconditional distribution $p(\mathbf{x})$ as follows:
$\begin{split}p(\mathbf{x}|\mathbf{c})&=\mathcal{N}\left(x_{1},x_{2}|(c_{1},c_{2})^{T},I\right)\\ p(\mathbf{x})&=\mathcal{N}\left(x_{1},x_{2}|(0,0)^{T},5I\right).\end{split}$ (21)
where $c_{1}=-5,c_{2}=5$ in the experiment. In the mixture of Gaussian, we trained conditional and unconditional DDPMs to learn the distributions $p(\mathbf{x}|\mathbf{c})=\prod_{i=0}^{2}\mathcal{N}\left(\mathbf{x}|\boldsymbol{\mu}_{i},I\right)^{c_{i}}$ and $p(\mathbf{x})=\frac{1}{3}\sum_{\{\mathbf{e}_{i},i=0,1,2\}}p(\mathbf{x}|\mathbf{e}_{i})$, where $\mathbf{c}=(c_{0},c_{1},c_{2})^{T}$ represents a three dimensional one-hot vector, $\boldsymbol{\mu}_{0}=(-1,-1/\sqrt{3})^{T}$, $\boldsymbol{\mu}_{1}=(1,-1/\sqrt{3})^{T}$, $\boldsymbol{\mu}_{2}=(0,\sqrt{3}-1/\sqrt{3})^{T}$, $\{\mathbf{e}_{i},i=0,1,2\}$ represents all one-hot vectors in three dimension space.
@@
In both experiments, we evaluated four different sampling methods: SDE (35) (1000 steps), probabilistic ODE (Song et al., 2020b) (1000 steps), DDIM (Song et al., 2020a) (20 steps), and DPM++2M (Lu et al., 2022) (20 steps). The projection operator $\mathbf{P}$ for characteristic guidance was set to identity. Samples were compared against the theoretical reference using Kullback–Leibler (KL) divergence.
@@
In the conditional Gaussian scenario, Fig.2 shows that characteristic guidance achieves better KL divergence than classifier-free guidance for every sampling method. It’s worth noting that classifier-free guidance suffers from severe bias and catastrophic loss of diversity when ODE based sampling methods (ODE, DDIM, and DPM++2M) are used. Contrarily, characteristic guidance always yields correct sampling with KL divergence better than classifier-free guidance’s best SDE samples. In the Mixture of Gaussian scenario, Fig.3 shows that characteristic guidance again showed less bias and better diversity. However, the overlap regions in the mixture of Gaussian scenario led to less clear component boundaries. This is because the harmonic ansatz is not valid in the overlap regions. Besides, minor artifacts appear because iterating on the trained neural network $\boldsymbol{\epsilon}$ brings in approximation errors. Overall, characteristic guidance outperformed classifier-free guidance in reducing bias, increasing diversity, and narrowing the performance gap between ODE and SDE sampling methods. This was consistently observed in cases where the harmonic ansatz holds exactly or approximately. More results could be found in Fig.8 and Fig.9.
@@ 6.2 Cooling of the Magnet: Where the Harmonic Ansatz Not Supposed to Hold
Figure 4: Comparison between CH (ours) and CF guided DDPM on simulating magnet cooling. The images in the upper and middle left (blue and red lattices) depict samples from DDPMs, while the histograms illustrate the distribution of samples’ mean value (magnetization) across different temperatures. The contours corresponds to the theoretical reference distribution of magnetization. The characteristic guidance has better NLL and is more capable in capturing peak separation of sample magnetization.
Figure 5: Left: CIFAR-10 aircraft images generated via DDPM highlight the difference between Classifier-free Guidance (CF) and Characteristic Guidance (CH) across various guidance scales ($\omega$). The CF-guided images tend to have dull or even white backgrounds at higher $\omega$, whereas the CH-guided images creates more vibrant scenes with skies and clouds. Right: In ImageNet 256, volcano samples generated using latent diffusion models with CF and CH Guidance without cherry-picking. CF images show color cast and underexposure at higher $\omega$, while CH images maintain consistent color and exposure, better highlighting volcanic features, such as smoke and lava. For these visual comparisons, consistent initial noise ensures that both CF and CH guided images maintain similar contexts at lower $\omega$ values.
In this section, we explore a simulation of the cooling of a magnet, comparing our method’s performance with that of classifier-free guidance. This simulation is not just a representation of a complex real-world physical phenomenon but also a test of our characteristic guidance on a case where the harmonic ansatz does not hold.
We focus on the cooling around the Curie temperature, a critical point where a paramagnetic material undergoes a phase transition to permanent magnetism. This transition, driven by 3rd order terms in the score function, challenges the harmonic ansatz. Despite this, our method successfully captures the most important phenomenon of this transition: the separation of two distinct peaks below the Curie temperature, which classifier-free guidance fails to reproduce. These results emphasize the robustness and effectiveness of our method in modeling intricate physical processes, even in scenarios where the harmonic ansatz does not hold.
@@
The probability of observing a particular configuration of $\phi$, according to statistical physics, is proportional to the Boltzmann distribution
@@
where $\beta H(\phi;T)$ is Hamiltonian at temperature $T$, assigning an energy to each of possible $\phi$. At temperature near the Curie temperature $T_{c}$, a Landau-Ginzburg model (Kardar, 2007) of magnet uses a Hamiltonian (67) (Appendix H) with 4-th order terms. Such terms are the key to characterise phase transition of magnet, but its gradient (score function) is a third order term that does not respect the harmonic ansatz.
@@
where $\omega$ is the guidance scale, $T_{1}>T_{0}$ are two distinct temperature values. This means if we know the distribution of our magnet at two distinct temperatures $T_{0}$ and $T_{1}$, we know its distribution at any temperature cooler than $T_{1}$ by adjusting the guidance scale $\omega$.
@@
We train a conditional DDPM simulating the cooling of a magnet with the Curie temperature $T_{c}=200$. The model is trained at two distinct temperatures $p(\phi|T_{0}=201)$ and $p(\phi|T_{1}=200)$, corresponds to $\omega=-1$ and $\omega=0$. Samples at lower temperatures $T=196$ ($\omega=4$) are later obtained by adjusting the guidance scale. Four distinct sampling methods, identical to those discussed in Sec.6.1, are evaluated. The projection operator $\mathbf{P}$ for characteristic guidance is set to be the channal-wise mean.
Fig.4 demonstrates the phase transition of mean magnetization, which presents the histograms of the mean value of $\phi$ sampled from the trained DDPM (8192 samples). Above the Curie temperature $T_{c}=200$, the histogram of the mean magnetization has one peak centered at 0. Below the Curie temperature, the histogram of the mean magnetization has two peaks with non-zero centers. More details are available in Fig.10. The change in the number of peaks represents a phase change from non-magnet to a permanent magnet.
@@
Characteristic guidance outperforms classifier-free guidance in sample quality below the Curie temperature, showcasing better negative log-likelihood (NLL) and effectively capturing distinct peaks in magnetization of samples. In contrast, classifier-free guidance struggles to produce distinct peaks. Despite its slightly biased peak positions compared with theoretical expectations, characteristic guidance still demonstrates superior performance in describing distinct peaks and modeling phase transitions.
6.3 CIFAR-10: Natural Image Generation